OpenStudy (anonymous):

what is the y-intercept of y=-2x-7 and how do u get it

OpenStudy (anonymous):

The y-intercept is -7 and it is gotten from the value of "b" in the slope-intercept form of a line: y = mx + b Your equation is already in that form, so you just match up the b with the "-7".

OpenStudy (anonymous):

put x=0. it will be your answer...

OpenStudy (anonymous):

-2x-7=-9x

OpenStudy (anonymous):

Incidentally, that form also gives you your slope as "m" or -2

OpenStudy (anonymous):

i have to graph it though

OpenStudy (anonymous):

so y=-9/x... i think

OpenStudy (anonymous):

Looks like:

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @kaitlyn_casteel